OpenStudy (anonymous):

Of the 30 students in the first period algebra I class, 18 are girls. If this class is representative of all classes, how many students in a school of 500 are girls?

OpenStudy (saifoo.khan):

use the Ratio method.

OpenStudy (saifoo.khan):

30 : 18 500 : x

hero (hero):

18/30 = x/500

OpenStudy (saifoo.khan):

300 girls

OpenStudy (anonymous):

I VOTE HERO AND CALLISTO

OpenStudy (anonymous):

WELL DONE

OpenStudy (callisto):

ammendment: 18/30 = x/500 x=300

OpenStudy (saifoo.khan):

Im also correct. lol

OpenStudy (anonymous):

can i haz medal?...and chezburger?

hero (hero):

amendment*

OpenStudy (anonymous):

fine i <3 u 2

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

now give me pwease